OpenStudy (anonymous):

Please help with his calc equation. Write an equation of the tangent function with period pi/4, phase shift pi, and vertical shift 1.

OpenStudy (ranga):

tan(x) has a period of pi. tan(ax) has a period of pi/a So to have a period of pi/4, pi/a = pi/4. So a = 4 tan(4x) has a period of pi/4 Phase shift of pi units to the right means, subtract pi from x. tan(4(x - pi)) vertical shift of + 1 means add 1 to the function. So the function f(x) = tan(4x - 4pi) + 1

OpenStudy (ranga):

And since tan repeats itself every pi radians, subtracting the 4pi will have no effect and therefore can be taken out. tan(4x - 4pi) + 1 = tan(4x) + 1